# MESA: Matching Everything by Segmenting Anything

### 30 Jan 2024 v1

@@ 4.4 Global Matching Energy Minimization
$E_{G}(v^{1}_{{h}})$ $=\frac{1}{Z}(\mu\cdot E_{self}(v^{1}_{{h}})+\alpha\cdot E_{parent}(v^{1}_{{h}})$ (13)
$+\beta\cdot E_{children}(v^{1}_{{h}})+\gamma\cdot E_{neighbour}(v^{1}_{{h}}))$
where $\mu,\alpha,\beta$ and $\gamma$ are weights to balance the terms; $Z$ is the partition function. The $E_{self}(v^{1}_{{h}})$ is the energy related to matching probability between $v^{0}_{src}$ and $v^{1}_{{h}}$.
$E_{self}(v^{1}_{{h}})=|\!|1-Sim_{a^{0}_{src},a^{1}_{h}}|\!|_{1}$ (14)
The $E_{parent}(v^{1}_{{h}})$ is the energy related to matching probability between the parent node pairs of $v^{0}_{src}$ and $v^{1}_{{h}}$.
$\centering\begin{gathered}E_{parent}(v^{1}_{{h}})=\\ \inf\{|\!|1-Sim_{a^{0}_{u},a^{1}_{r}}|\!|_{1}\big{|}\small u\in p^{0}(src),r\in p^{1}(h)\}\end{gathered}\@add@centering$ (15)
where $p^{i}(j)$ is the index set of parent nodes of $v^{i}_{j}$ in $\mathcal{G}_{i}$. This energy is the minimum matching energy among all parent node pairs of $v^{1}_{{h}}$ and $v^{0}_{src}$. Same as the $E_{parent}$, the $E_{children}$ and $E_{neighbour}$ are the energy parts of children and neighbour node matching pairs. Afterwards, the best area match $v^{1}_{h^{*}}$ in the set can be found by minimize $E_{G}$.
$h^{*}=\arg\min_{h\in{\mathcal{H}}}E_{G}(v^{1}_{h})$ (16)
@@ 5 Experiments
SEEM [51]+SGAM [49$\dagger$ 52.96 49.39 29.88 12.32 2.66
@@ 5.1 Area Matching
Results. In Sec. 5, our MESA outperforms SGAM by a large margin ($60.5967.98~{}v.s.~{}$ 60.59 in AOR). The poor area matching performance of SGAM is mainly caused by the inaccurate semantic labeling of SEEM, which is avoided in MESA. MESA can also achieve more areas than SGAM, favouring downstream tasks by more point matches. Although GAM brings significant improvement for SGAM, it achieves less enhancement on MESA, because more precise area matching is obtained by MESA. We empirically adopt GAM w/ $\phi=1.0$ for MESA in the following experiments.
@@ 5.2 Indoor Pose Estimation
Experimental setup. The parameters of MESA are the same as our area matching experiment in Sec. 5.1. We adopt the A2PM framework [49] and combine four different state-of-the-art point matchers with our method, i.e. semi-dense LoFTR [38], ASpan [8], QuadT [40] and dense DKM [16], to demonstrate the performance improvement and compare with recent matching methods [18, 6, 20]. The area images are resized to the default input size of combined baseline for point matching. Following [38, 8], the pose estimation AUC at various thresholds are used as metrics.
Results. As we can see in Tab. 2, MESA significantly increases the pose accuracy for all baselines. Due to the presence of more severe matching redundancy in indoor scenes, our approach gains more noticeable enhancement, e.g. ${+15.26\%}$ in AUC@$5^{\circ}$ for QuadT, compared to outdoor scenes in Tab. 3. The improvement obtained by MESA for semi-dense matchers are comparable to those of SGAM [49], because the coarse-to-fine pipeline of these methods makes them insensitive to AOR. However, MESA brings a larger improvement for DKM [16] than SGAM ($13.61\%~{}v.s.~{}4.12\%$) and boosts the precision to a new state-of-the-art, proving the precise area matching by MESA benefits the performance of dense matching.
@@
5.3 Outdoor Pose Estimation
Experimental setup. The parameter $T_{E_{max}}$ is set as $0.3$. MESA is also combined with four baselines to compare with recent methods [45, 28, 17]. The cropped areas are resized to the default input size of combined point matcher. The pose estimation AUC is adopted as the metric.
Results. In Tab. 3, MESA achieves impressive precision improvement for semi-dense baselines, up to $+7.58\%$ in AUC@$5^{\circ}$. The improvement for the dense method is not as notable as that for semi-dense methods, because lots of repetitiveness in MegaDepth leads to hard area matching, decreasing the accuracy of DKM. However, MESA_DKM does set a new state-of-the-art for outdoor pose estimation.
@@ 5.4 Visual Odometry
Experimental setup. The parameter $T_{E_{max}}$ is set as $0.25$. The area image size is $640\times 480$. Our method is combined with four baselines [8, 40, 38, 16] to demonstrate its effectiveness. All baselines are trained on ScanNet [11]. Following [24], we report the relative pose errors (RPE), including the rotational error ($R_{err}$) and translation error ($t_{err}$).

### 08 Apr 2024 v2

@@ 4.4 Global Matching Energy Minimization
[…]
where $\mu,\alpha,\beta$ and $\gamma$ are weights to balance the terms; $Z$ is the partition function. The $E_{self}(v^{1}_{{h}})$ is the energy related to matching probability between $v^{0}_{src}$ and $v^{1}_{{h}}$:
$E_{self}(v^{1}_{{h}})=|1-Sim_{a^{0}_{src},a^{1}_{h}}|.$ (12)
The $E_{parent}(v^{1}_{{h}})$ is the energy related to matching probability between the parent node pairs of $v^{0}_{src}$ and $v^{1}_{{h}}$:
$\centering\begin{gathered}E_{parent}(v^{1}_{{h}})=\\ \min\{|1-Sim_{a^{0}_{u},a^{1}_{r}}|\big{|}\small u\in p^{0}(src),r\in p^{1}(h)\},\end{gathered}\@add@centering$ (13)
where $p^{i}(j)$ is the index set of parent nodes of $v^{i}_{j}$ in $\mathcal{G}_{i}$. This energy is the minimum matching energy among all parent node pairs of $v^{1}_{{h}}$ and $v^{0}_{src}$. Same as the $E_{parent}$, the $E_{children}$ and $E_{neighbour}$ are the energy parts of children and neighbour node matching pairs. Afterwards, the best area match $v^{1}_{h^{*}}$ in the set can be found by minimize $E_{G}$:
$h^{*}=\arg\min_{h\in{\mathcal{H}}}E_{G}(v^{1}_{h}).$ (14)
@@ 5 Experiments
SEEM [54]+SGAM [52$\dagger$ 52.96 49.39 29.88 12.32 2.66
@@ 5.1 Area Matching
Results. In Sec. 5, our MESA outperforms SGAM by a large margin ($60.5967.98~{}vs.~{}$ 60.59 in AOR). The poor area matching performance of SGAM is mainly caused by the inaccurate semantic labeling of SEEM, which is avoided in MESA. MESA can also achieve more areas than SGAM, favouring downstream tasks by more point matches. Although GAM brings significant improvement for SGAM, it achieves less enhancement on MESA, because more precise area matching is obtained by MESA. We empirically adopt GAM w/ $\phi=1.0$ for MESA in the following experiments.
@@ 5.2 Indoor Pose Estimation
Experimental setup. The parameters of MESA are the same as our area matching experiment in Sec. 5.1. We adopt the A2PM framework [52] and combine four fine-tuned point matchers with our method, i.e. LoFTR [40], ASpan [8], QuadT [43] and DKM [17], to demonstrate our performance and compare with recent methods [19, 6, 21]. The cropped area images are resized to be square with the same side length as the long side of the default size of point matchers. Following [40], the pose estimation AUC at various thresholds are used as metrics. More implementation details can be found in our Supp. Mat.
Results. As we can see in Tab. 2, MESA significantly increases the pose accuracy for all baselines. Due to the presence of more severe matching redundancy in indoor scenes, our approach gains more noticeable enhancement, e.g. ${+15.26\%}$ in AUC@$5^{\circ}$ for QuadT, compared to outdoor scenes in Tab. 3. The improvement obtained by MESA for semi-dense matchers are comparable to those of SGAM [52], because the coarse-to-fine pipeline of these methods makes them insensitive to AOR. However, MESA brings a larger improvement for DKM [17] than SGAM ($13.61\%~{}vs.~{}4.12\%$) and boosts the precision to a new state-of-the-art, proving the precise area matching by MESA benefits the performance of dense matching.
5.3 Outdoor Pose Estimation
Experimental setup. The parameter $T_{E_{max}}$ is set as $0.3$. MESA is combined with four baselines to compare with recent methods [48, 29, 18]. We also compare with another covisible-area matching method, OETR [9], combined with our post-processing. The area images are cropped and resized as input of point matcher, same as the indoor experiment. The pose estimation AUC is adopted as the metric.
@@ 5.3 Outdoor Pose Estimation
Results. In Tab. 3, MESA achieves impressive precision improvement for semi-dense baselines, up to $+7.58\%$ in AUC@$5^{\circ}$. The improvement for the dense method is not as notable as that for semi-dense methods, because lots of repetitiveness in MegaDepth leads to hard area matching, decreasing the accuracy of DKM. However, MESA_DKM does set a new state-of-the-art for outdoor pose estimation. MESA also achieves better results than OETR by finding more detailed area matches, whereas OETR solely estimates covisible areas.
@@ 5.4 Visual Odometry
Experimental setup. The parameter $T_{E_{max}}$ is set as $0.25$. The area image size is $480\times 480$. Our method is combined with four baselines [8, 43, 40, 17] to demonstrate its effectiveness. All baselines are trained on ScanNet [11]. Following [25], we report the relative pose errors (RPE), including the rotational error ($R_{err}$) and translation error ($t_{err}$).
@@ 5.5 Ablation Study
Supplementary Material
In the following, we provide additional details about the proposed area matching method, MESA. Sec. 7 describes specific implementation details of our method. Sec. 8 analyzes the computation complexities of main components in MESA. Sec. 9 gives more insights and ablation studies, that motivates our design. Sec. 10 states the limitation of MESA and our future work.
7 Implementation Details
In this section, we provide sufficient details about the implementation of MESA for reproduction, including detailed operations in Graph Completion (Sec. 7.1), training details about our Learning Area Similarity (Sec. 7.2) and complete parameter settings in Graphical Area Matching (Sec. 7.3).
7.1 Graph Completion
The detailed algorithm for our graph completion is depicted in Algorithm 1, which takes initial AG ($\mathcal{G}_{ini}$) as input and outputs the final AG ($\mathcal{G}$) with scale hierarchy. Additionally, we describe the area cluster and two main area operations in the algorithm as follows.
Area Cluster. For orphan nodes in each level, we cluster them based on their area centers to decide which operation will be performed on them. We use the k-means algorithm with elbow method [41] to determine the cluster number. The candidate cluster number is set as $\{1,\dots,n\}$, where $n$ is the number of orphan nodes in the current level. This algorithm is fed with area centers and outputs labeled ones.
Area Fusion and Expansion. Area fusion and expansion are key operations in our graph completion algorithm. Specifically, area fusion is to find the largest outer rectangle of the two areas as the new area, as depicted in Fig. 6 (a). Due to the careful threshold settings of our area level, the fused area size will exceed current level and be awaited for subsequent operations. On the other hand, the expansion operation is to expand the area to the next level size (Fig. 6 (b)). In particular, suppose the lower bound of size for the next level is $s^{2}$, if both of the area width and height are smaller than $s$, we expand the height and width of the area to $s$, keeping the area center fixed. Otherwise if area width $w\geq s$, we let the area height $h=s^{2}/w$, keeping the area center fixed, and vice versa. The above operations are performed when the expanded area is inside the image. On the other hand, if the expanded area is outside the image, the area center will be moved as shown in Fig. 6 (b).
7.2 Learning Area Similarity.
We propose the learning area similarity model for area similarity calculation, which is the basis of our graphical area matching. In this section, we describe the training protocol of this model, including supervision and training details.
Supervision. We generate regular area images from both indoor and outdoor datasets [11, 27] as training data through the proposed method in Sec. 3. Then the area pairs with more than $30\%$ overlap are collected. For each image patch $p_{i}$ in these pairs, its ground truth activity $\sigma_{i}^{gt}$ is set as $1$, if more than $60\%$ pixels in it have correspondences in the other area image, and 00 otherwise. As our classification formulation, we use the binary cross entropy ($BCE$) of each patch classification to form the loss function of area similarity calculation ($L_{asc}$).
$L_{asc}=\frac{1}{Z}\sum_{i}^{Z}BCE(\sigma_{i}^{gt},\sigma_{i})$ (16)
Based on this loss, our network can learn to achieve the similarity between two area images.
Input: $\mathcal{G}_{ini}=\langle\mathcal{V}_{ini},\mathcal{E}_{ini}\rangle$
Output: $\mathcal{G}=\langle\mathcal{V},\mathcal{E}\rangle$
1 for $l$ in $[0,L-1]$ do
2       initial orphan node set $\mathcal{O}=\varnothing$;
3       for $v_{i}\in\{v_{i}|l_{a_{i}}=l\}$ do
4             if $v_{i}$ has no parent then
5                   add $v_{i}$ into $\mathcal{O}$;
6
7
8      cluster the nodes in $\mathcal{O}$ based on their area centers;
9       for each node cluster $\mathcal{C}_{h}=\{v_{k}\}_{k=0}^{C}$ do
10             if $C\geq 2$ then
11                   for each $v_{k}\in\mathcal{C}_{h}$ do
12                         if $v_{k}$ has not been fused then
13                               fuse area $a_{k}$ with its nearest neighbor $a^{n}|v^{n}\in\mathcal{C}_{h}$: $a^{f}=F({a}_{k},a^{n})$;
14                               generate higher level node $v^{f}$ for $a^{f}$;
15                               add $v^{f}$ into $\mathcal{V}_{ini}$;
16                               form edges by Link Prediction: $\{e_{h}\}_{h}=LP(v^{f},\mathcal{V}_{ini})$;
17                               add $\{e_{h}\}_{h}$ into $\mathcal{E}_{ini}$;
18
19
20
21            else
22                   Update the single node $v_{0}$: $v^{u}_{0}=Up(v_{0})$;
23                   construct edges: $\{e_{j}\}_{j}=LP(v^{u}_{0},\mathcal{V}_{ini})$;
24                   add $\{e_{j}\}_{j}$ into $\mathcal{E}_{ini}$;
25
26
27
28$\mathcal{E}=\mathcal{E}_{ini}$ ;
29 $\mathcal{V}=\mathcal{V}_{ini}$ ;
30 output the updated AG: $\mathcal{G}=\langle\mathcal{V},\mathcal{E}\rangle$;
Algorithm 1 Graph Completion
Figure 6: The Area Fusion and Area Expansion. (a) Area fusion is to achieve the smallest area (C) containing the input areas (A and B). (b) Generally, area expansion is to fix the original area center and expand its size to the smallest size of the next level. When the original area is too close to the image boundary, we will move the area center to keep the expanded area inside the image.
Training Details. Following previous point matching work [40], the proposed learning similarity module is trained utilizing MegaDepth [27] and ScanNet [11] dataset. For each image pair sampled in the datasets in every epoch following [40], we collect up to 5 area image pairs for training. Fortunately, there is no need to train this model from scratch, thanks to the similar objective between the coarse point matching and the area similarity calculation, both of which aims at patch-level similarities. Therefore, we adopt the pretrained weights of coarse level feature operation in ASpan [8] on our network. With the modified output head, our network is fine-tuned using AdamW [30] on 2 NVIDIA RTX 4090 GPUs. We use the same learning rate and batch size settings as ASpan and LoFTR [40] during training. This model can converge within 3 epochs.
7.3 Graphical Area Matching
In this section, we describe some implementation details of our graphical area matching. First, we illustrate the complete forward process of area matching. Then, we specific the parameter settings we recommend and utilized in our experiments. Next, the details about area image cropping is presented, which is cropping area images from original images to serve as input images for the combined point matcher. The post-processing of point matching within areas is also illustrated.
Area Matching Process. Given the image pair ($I_{0},I_{1}$) and their area graphs ($\mathcal{G}^{0},\mathcal{G}^{1}$), we first collect area nodes with specific size level $l_{a^{*}}$ from $\mathcal{G}^{0}$ as the source nodes, which have proper sizes for subsequent point matching. Then, area matches are found for these nodes from $\mathcal{G}^{1}$ by our method. Afterwards, we exchange the two images to repeat the above operations. The final area matches are the common matching results.
Parameter Settings. We describe the common parameters for different scenes here. During the AG construction, the input image are resized to $640\times 480$. The aspect ratio threshold $T_{r}=4$ and minimal size threshold is $T_{s}=80^{2}$. The number of area size threshold is $4$ and specific $TL_{i}$s are $\mathsf{80^{2},130^{2},256^{2},390^{2},560^{2}}$. The $\delta_{l}$ is $0.1$ and $\delta_{h}$ is $0.8$. In graphical area matching, the $\lambda$ in Eq. 5 is $0.1$. The area similarity threshold $T_{as}=0.05$. The energy balance weights ($\mu,\alpha,\beta,\gamma$) in Eq. 11 are $\mathsf{4,2,2,2}$. The specific area level $l_{a^{*}}$ for point matching is $1$. The $T_{Er}$ in Eq. 15 is $0.1$. Other parameters are specified for different scenes as described in our paper.
Area Image Cropping and Point Matching. Following the A2PM framework [52], the next step after area matching is to perform point matching inside area matches. Thus, these area matches need to be achieved through image cropping. At the same time, one of the benefits of A2PM is the input image with high resolution that can be provided to the point matcher. Therefore, the cropping is performed on original images with the highest resolution. A straightforward cropping approach is to crop areas along the exact bounding box and then resize them to default input size of point matchers. However, as area sizes can be quite different from default input sizes of point matchers, direct cropping and resize introduce severe distortion for point matching, resulting in decreased matching precision (cf. Sec. 9.1). In order to address this issue, we propose to crop area image by considering the aspect ratio. To be specific, we force the cropped area image to possess the same aspect ratio as the input size of the point matcher, while trying to keep the area center unchanged. If the area respect ratio ($W_{a}/H_{a}$) is larger than the input aspect ratio ($W_{i}/H_{i}$), we fix the width $W_{A}$ of area image and expand the height $H_{a}$ to $(W_{a}\times H_{i})/W_{i}$. Otherwise, we fix the height and expand the width to $(H_{a}\times W_{i})/H_{i}$. Moreover, as solid feature points often cluster in the boundaries of objects, we set a spread ratio ($r_{s}=1.2$) to slightly increase the cropping size, allowing for more precise correspondences on the boundaries. Finally, if the cropped area exceeds the original image, we will move its center to keep it inside the image, similar to our Area Expansion in Fig. 6.
For point matching inside area matches, in practice, we empirically set the input size of point matchers to a square, i.e. the input aspect ratio is $1$, as it can lead to statistically the smallest area size adjustment. However, as the learning models are sensitive to training size and scale (the area input has different scale with original image), especially the Transformer [13], the performance of baselines are decreased, as we shown in Sec. 9.1. Therefore, we fine-tune the baselines under the square sizes with ground truth area matches generated by our method, e.g. $640\times 640$ area matches in indoor scenes, to achieve similar accuracy with original models, before combined with MESA.
Post-processing. In SGAM [52], Global Match Collection (GMC) is proposed to collect precise point matches through globally entire-image matching, which significantly improve the matching precision [52]. Therefore, we adopt this module in our method as well and set the occupancy ratio as $0.6$. However, as SAM enables areas throughout the image, MESA usually achieves enough area matches to cover the whole overlap between images. Thus, this module is only activated in few cases, but also helpful to precise matching. As we mentioned before, we also adopt GAM [52] as the post-processing of area matching.
8 Computation Complexity
Here, we analyze the computation complexity of proposed graphical area matching.
8.1 Area Similarity Calculation
Firstly, area similarity calculation is performed to achieve the required node energies in the graph, serving as the prerequisite of our graphical area matching. Suppose we have two AGs, $\mathcal{G}^{0}$ and $\mathcal{G}^{1}$, for the input image pair, $\mathcal{G}^{0}$ gets $N$ nodes ($|\mathcal{V}^{0}|=N$) and $\mathcal{G}^{1}$ gets $M$ nodes ($|\mathcal{V}^{1}|=M$). Therefore, the dense graph energy calculation needs $M\times N$ times similarity calculation. However, owing to the similarity conditional independence of ABN (Sec. 4.3), the actual number ($M^{\prime}\times N^{\prime}$) of similarity calculation is smaller than $M\times N$, as $N^{\prime}<N$. Nevertheless, directly setting children pair similarities as 00 is too rough (Eq. 10), as large scale differences also leads to near-zero similarity between areas. In practise, we only set the related similarities of next level children as 00 for area matching accuracy and the efficiency from ABN is still helpful to our approach. Moreover, we only care about the similarities between source nodes in $\mathcal{G}^{0}$ and other nodes in $\mathcal{G}^{1}$, because we collect source nodes with specific level from $\mathcal{G}^{0}$ to match, e.g., usually $3\sim 4$ areas in indoor scene and less in outdoor scene. Therefore, we have $M^{\prime}<M$. Similarly, in the case of duality, i.e., collecting source nodes from $\mathcal{G}^{1}$ to match, we only need to perform a few supplementary calculations, as similarities are symmetric and reusable. Thus, the real computation complexity of area similarity computation is $O(M^{\prime}\times N^{\prime})$, where $M^{\prime}\times N^{\prime}<M\times N$.
8.2 Edge Energy Calculation
Except the node energy calculation, the edge energy is also needed to be determined for Graph Cut. The computation complexity of edge energy calculation is related to edge number of $\mathcal{G}^{0}$ and $\mathcal{G}^{1}$. Assume $|\mathcal{E}_{0}|=E$ and $|\mathcal{E}_{1}|=K$, the specific computation complexity is $O(E+K)$.
8.3 Global Energy Minimization
In our global energy minimization for area matching refinement, the matching energy of parent, children and neighbour pairs all need to be calculated. Taking parent matching energy for example, we derive its computation complexity as follows. Suppose $n$ nodes are achieved as match candidates through Graph Cut and each node gets $Q_{i},\;i\in(0,n]$ parent nodes, there are $Q_{i}\times V$ node similarities need to be accessed (as the similarity calculation is finished), where $V$ is the parent node number of the source node. Hence, the total computation complexity for parent matching energy in global energy minimization is $O(\sum_{i}^{n}Q_{i}\times V)$. The children matching energy and neighbour matching energy are similar. As $n$ is the number of node after Graph Cut, it is small in most cases, e.g., usually $<3$ area nodes. Moreover, the number of parent nodes (or children, neighbour nodes) is also limited. Therefore, the computation complexity for global energy minimization is acceptable in practise.
Table 6: Ablation study of area image size. We investigate the performance impact of three different input image sizes for both semi-dense and dense methods, along with our MESA combined with them. For point matchers, the input image size is the size of resized original image. For our method, the input image size is the area image size. The pose estimation AUC@$5^{\circ}/10^{\circ}/20^{\circ}$ are reported for evaluation.
Input Image Size Semi-Dense Matcher AUC@5 $\uparrow$ AUC@10 $\uparrow$ AUC@20 $\uparrow$ Dense Matcher AUC@5 $\uparrow$ AUC@10 $\uparrow$ AUC@20 $\uparrow$
$640\times 640$ ASpan 26.33 45.98 62.12 DKM 28.63 50.84 68.97
MESA_ASpan 27.51${}_{+4.48\%}$ 47.47${}_{+3.24\%}$ 65.04${}_{+4.70\%}$ MESA_DKM 33.42${}_{+16.73\%}$ 55.04${}_{+8.26\%}$ 71.98${}_{+4.36\%}$
[…]
MESA_ASpan 24.01${}_{+12.30\%}$ 43.32${}_{+6.88\%}$ 60.99${}_{+3.83\%}$ MESA_DKM 33.00${}_{+14.38\%}$ 54.04${}_{+7.95\%}$ 71.02${}_{+4.13\%}$
[…]
MESA_ASpan 9.60${}_{+7.26\%}$ 22.17${}_{+4.67\%}$ 38.70${}_{+2.71\%}$ MESA_DKM 31.43${}_{+14.08\%}$ 52.56${}_{+9.05\%}$ 69.99${}_{+6.11\%}$
Method Cropping Approach AUC@5 $\uparrow$ AUC@10 $\uparrow$ AUC@20 $\uparrow$
MESA_ASpan OAR 24.67 43.72 61.29
ARPM 27.51 47.47 65.04
[…]
ARPM 33.42 55.04 71.98
Table 7: Ablation study of area image cropping. Two different image cropping methods are compared for the proposed MESA. Both semi-dense and dense point matchers are combined for evaluation. We report the pose estimation AUC@$5^{\circ}/10^{\circ}/20^{\circ}$ and the best results of two series are highlighted respectively.
9 Additional Ablation Study
In this section, we examine the performance impact of more components in MESA, including the input image size (Sec. 9.1), image cropping approach (Sec. 9.2) and energy parameter setting (Sec. 9.3).
9.1 Ablation Study on Input Image Size
[…]
9.2 Ablation Study on Image Cropping
The image cropping is a trivial yet important operation for the A2PM framework, as different cropping approaches lead to different image resolutions and distortions. Here, we construct experiments to investigate the impact of different cropping approaches. To be specific, we compare two different cropping methods descried in Sec. 7.3: 1) the direct cropping method (OAR), which crops with Original Aspect Ratios of areas; 2) the cropping method with the Aspect Ratio of Point Matcher (ARPM), which first expands the area to correspond with the aspect ratio of the point matcher input and then crop areas. The experiment is conducted on ScanNet1500 [11] benchmark. We combine MESA with both semi-dense (ASpan) and dense (DKM) point matchers for complete comparison. Results are summarized in Tab. 8.3. As we can seen that the ARPM cropping approach outperforms the OAR approach with a large margin for both MESA_ASpan and MESA_DKM, proving its superiority due to high resolution and less distortion. Therefore, we adopt the ARPM approach for area image cropping in MESA.
Table 8: Ablation study of global energy parameters. We compare different parameter settings for global energy refinement in MESA_ASpan and report the area matching performance, area number per image (AreaNum), and the pose estimation performance. Results are highlighted as first, second and third.
$E_{G}$ Parameters $T_{E_{max}}$ AOR $\uparrow$ AMP@$0.6$uparrow Pose AUC@$5^{\circ}$uparrow AreaNum $\uparrow$
$\mu=5,\;\alpha=2,$ $\beta=2,\;\gamma=1$ 0.35 61.76 65.54 23.57 4.69
0.25 63.91 71.13 22.41 3.47
0.15 60.44 62.57 21.46 3.27
[…]
0.25 64.94 72.24 24.01 4.62
0.15 61.74 65.50 23.55 3.86
[…]
0.25 62.32 66.54 23.56 2.92
0.15 60.32 64.38 22.37 2.77
Method LoFTR [40] PATS [21] DKM [17] COTR [20] MESA (Ours)
Time(ms)/img 181.5 450.4 1318.5 18975.3 3092.4
Table 9: Time consumption comparison. All methods run on $640\times 480$ images.
Table 8: Ablation study of global energy parameters. We compare different parameter settings for global energy refinement in MESA_ASpan and report the area matching performance, area number per image (AreaNum), and the pose estimation performance. Results are highlighted as first, second and third.
Table 6: Ablation study of area image size. We investigate the performance impact of three different input image sizes for both semi-dense and dense methods, along with our MESA combined with them. For point matchers, the input image size is the size of resized original image. For our method, the input image size is the area image size. The pose estimation AUC@$5^{\circ}/10^{\circ}/20^{\circ}$ are reported for evaluation.